OpenStudy (anonymous):

How do i find the distance between these points? (2,2,-9) and (-1,6,0)

OpenStudy (anonymous):

I understand how to find the distance when there is only (x,y) but when its (x,y,z) im lost.

OpenStudy (jit4won):

use the same formula Just add (z2-z1)^2 with them

OpenStudy (jit4won):

Getting it ?

OpenStudy (anonymous):

Can you show me the equation with that added?

OpenStudy (jit4won):

Check it out

OpenStudy (anonymous):

\[\sqrt{(-1-2)^2+(6-2)^2+(0+9)^2}\] \[\sqrt{(-3)^2+(4)^2+(9)^2}\] \[\sqrt{9+16+81}\] \[\sqrt{106}\]

OpenStudy (anonymous):

How come you add the last one? and subtract all the others?